Pearl Corp. is expected to have an EBIT of $2,900,000 next year. Depreciation, the increase in net working capital, and capital spending are expected to be $160,000, $130,000, and $170,000, respectively. All are expected to grow at 19 percent per year for four years. The company currently has $15,000,000 in debt and 1,300,000 shares outstanding. At Year 5, you believe that the company's sales will be $23,840,000 and the appropriate price-sales ratio is 2.1. The company’s WACC is 8.5 percent and the tax rate is 21 percent.

What is the price per share of the company's stock?
Answer #1

Expected FCF, FCF1 = EBIT * (1 - Tax Rate) + Depreciation - Increase in NWC - Capital Spending
Expected FCF, FCF1 = $2,900,000 * (1 - 0.21) + $160,000 - $130,000 - $170,000
Expected FCF, FCF1 = $2,151,000

Growth rate for next 4 years is 19%

FCF2 = $2,151,000 * 1.19 = $2,559,690
FCF3 = $2,559,690 * 1.19 = $3,046,031
FCF4 = $3,046,031 * 1.19 = $3,624,777
FCF5 = $3,624,777 * 1.19 = $4,313,485

Price-Sales Ratio = Horizon Value of Firm / Sales
2.10 = Horizon Value of Firm / $23,840,000
Horizon Value of Firm = $50,064,000

WACC = 8.50%

Value of Firm = $2,151,000/1.085 + $2,559,690/1.085^2 + $3,046,031/1.085^3 + $3,624,777/1.085^4 + $4,313,485/1.085^5 + $50,064,000/1.085^5
Value of Firm = $45,320,636

Value of Equity = Value of Firm - Value of Debt
Value of Equity = $45,320,636 - $15,000,000
Value of Equity = $30,320,636

Price per share = Value of Equity / Number of shares outstanding
Price per share = $30,320,636 / 1,300,000
Price per share = $23.32